OpenStudy (delahruza):

Let G be a multiplicatively-written group with exactly two elements e and a, where e is the identity element. Prove that a^2=e. My proof: since there are only elements e and a, it means that a=a^(-1). e=a*a^(-1)=a*a=a^2 Logically that's probably fine, but can this be considered a correct proof?

OpenStudy (perl):

proof: if a = e, then e^2 = e*e = e , and you are done. it might be easier to show this with a cayley table

OpenStudy (perl):